OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Dolphin Tricks & Tours: Age Ticket Price Under 3 years free 3 to 12 years $35 Over 12 years $46 On a certain day, the company took 88 people on whale watching trips. There were 44 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (44 + x)35 + (88 - (44 + x)) ⋅ 46 = C (88 - x)35 + (88 - 44) ⋅ 46 = C (44 - x)35 + (88 - 44) ⋅ 46 = C (88 + x)35 + (88 - (44 + x)) ⋅ 46 = C

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

x+(between 3 and 12) = 44 right? so (between 3 and 12) = 44 - x then (number of people)*(cost) + (number of people)*(cost) = (total cost)

OpenStudy (anonymous):

88 * c = 88c

OpenStudy (anonymous):

im thinking either A or D

jigglypuff314 (jigglypuff314):

(between 3 and 12)*(cost) + (total people - under 12)*(cost of old) = total cost

OpenStudy (anonymous):

(3 to 12) * C + 88 - 44 * C = C Sorry if this is wrong, I can't think at all today...

jigglypuff314 (jigglypuff314):

it's alright, you have it right so far and I said (between 3 and 12) = 44 - x

OpenStudy (anonymous):

Oh wait! I think I get it, so it is C then, right?

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

Thank you! :)